# Personalized Reasoning: Just-in-time Personalization and Why LLMs Fail at It

### v1

@@ Abstract
Current large language model (LLM) development treats task-solving and preference-alignment as separate challenges, optimizing first for objective correctness, then for alignment to aggregated human preferences. This paradigm fails in human-facing applications where solving a problem correctly is insufficient if the response mismatches the user’s needs. This challenge intensifies in just-in-time scenarios where no prior user interaction history exists due to cold-start conditions or privacy constraints. LLMs need to identify what they don’t know about user preferences, strategically elicit preference values through questioning, then adapt their reasoning processes and responses accordingly—a complicated chain of cognitive processes which we term personalized reasoning. We introduce PrefDisco, an evaluation methodology that transforms static benchmarks into interactive personalization tasks using psychologically-grounded personas with sparse preferences. Our framework creates scenarios where identical questions require different reasoning chains depending on user context, as optimal explanation approaches vary by individual expertise and preferences while maintaining factual accuracy. Evaluation of 21 frontier models across 10 tasks reveals 29.0% of naive personalization attempts produce worse preference alignment than generic responses, yet generic responses also fail to serve individual user needs effectively. These findings suggest personalized reasoning requires dedicated development rather than emerging naturally. PrefDisco establishes personalized reasoning as a measurable research frontier and reveals fundamental limitations in current LLMs’ interactive capabilities, providing a foundation for developing systems that can adapt to individual users in education, healthcare, and technical domains where personalization is critical.
@@ 1 Introduction
Current large language model (LLM) development treats task-solving and preference alignment as sequential challenges: models are first optimized for objective correctness through instruction tuning or reinforcement learning (Longpre et al., 2023), then aligned to aggregated human preferences through reinforcement learning from human feedback (Ouyang et al., 2022; Rafailov et al., 2024). This paradigm fundamentally misaligns with human-AI interaction, where the task and the individual user are inseparable. For instance, a medical explanation benefits from clinical analogies for one trainee while another requires formal definitions, requiring different cognitive approaches to answer the same problem. When models provide identical responses regardless of user context, they fail to serve individual needs despite achieving high benchmark performance. This challenge intensifies in cold-start scenarios where no prior user interaction history exists due to privacy constraints or new user onboarding, requiring “just-in-time personalization” capabilities that current systems lack. Moreover, users often cannot articulate their specific needs or provide effective feedback about response misalignment (Liu et al., 2025), necessitating that LLMs proactively elicit this information rather than placing the cognitive burden on users.
@@
Existing personalization research inadequately addresses interactive reasoning scenarios. Personalization benchmarks such as PersoBench (Afzoon et al., 2024), PrefEval (Zhao et al., 2025), and PersonaMem (Jiang et al., 2025) focus on content recommendation or dialogue generation with static user profiles, treating personalization as applying predetermined preferences to fixed outputs rather than adapting the underlying reasoning approach. Interactive frameworks like MediQ (Li et al., 2024) demonstrate questioning capabilities but target clinical information-seeking without personalization objectives. Most critically, no existing work recognizes that effective personalization requires different reasoning processes for different users; current approaches assume reasoning processes remain constant while only presentation varies. We address this conceptual gap by introducing the first evaluation framework requiring models to discover user preferences and adapt their reasoning processes accordingly, recognizing that the cognitive steps needed to solve problems should themselves be user-dependent.
@@
Evaluation of 21 frontier models across 10 tasks reveals systematic failures in personalized reasoning capabilities. In 29.0% of cases, attempting personalization produces worse preference alignment than generic responses, yet the genetic responses still fail to address user needs. Models exhibit insufficient questioning, asking only 1.42 questions on average despite 5-turn allowances, and fail to identify relevant preference dimensions. Domain analyses show disparities among task types: mathematical reasoning suffers severe degradation under personalization constraints (3.5% accuracy loss) while social reasoning maintains robustness (3.1% gain). The (lack of) personalized reasoning capability in models has critical implications for educational applications, where misaligned explanations can impede learning by providing inappropriate cognitive scaffolding, and for healthcare and technical support domains where one-size-fits-all responses may lead to misunderstanding of complex procedures or safety-critical information. We make the following contributions:
  • We define personalized reasoning as a distinct capability requiring models to discover user preferences through strategic questioning and adapt their responses accordingly, distinguishing it from static persona consistency or content recommendation.
@@
In this section, we formalize personalized reasoning through a three-part decomposition. Section 2.1 introduces the notion of task-relevant attributes that define the space in which personalization can occur. Section 2.2 develops the representation of user preferences over these attributes and describes elicitation as a sequential decision process. Section 2.3 formalizes how models adapt their responses to the inferred preference profile and how alignment is evaluated jointly with correctness. This decomposition provides the foundation for the PrefDisco evaluation framework in Section 3.
@@ 3 PrefDisco Benchmark Construction
Traditional personalization assumes fixed preference profiles that apply universally across tasks. We reject this approach because psychological research demonstrates that individuals prioritize different attributes across contexts (Fischer & Schwartz, 2011). As formalized in §2.2, for each persona-problem pair $(p,i)$, we sample sparse preference profiles $\mathcal{P}_{p,i}=\{(\theta_{j},v_{j},w_{j})\}_{\theta_{j}\in\mathcal{F}(i)}$ where only context-relevant attributes are active. Further, we ground the preference sampling process on existing research in education, which states that frequently modeled student characteristics include knowledge level, misconceptions, cognitive features, affective features, and meta-cognitive features (Chrysafiadi et al., 2015).
@@
We implement passive user simulation inspired by (Li et al., 2024), where the user factually answers the model’s elicitation question without providing any extra details. The passive user faithfully represents the challenging scenario of preference discovery while minimizing confounding factors. The passive user type forces models to develop strategic questioning strategies rather than relying on user proactiveness and isolates models’ questioning capabilities from user communication style, providing controlled evaluation conditions. The 5-turn limit reflects realistic attention constraints in human-AI interaction while providing sufficient opportunity for effective preference discovery, as demonstrated by our correlation analysis between questioning volume and alignment quality.
@@ 4 Experiments
Our goal is to rigorously evaluate models’ ability for interactive preference discovery: engaging in dialogue to uncover hidden user requirements and adapt responses accordingly. To this end, our experiments combine three ingredients. First, we use diverse benchmarks spanning mathematical, scientific, and social reasoning to ensure domain-agnostic evaluation. Second, we introduce varied personas that encode heterogeneous user preferences, simulating realistic user variability. Third, we define controlled evaluation conditions that disentangle raw task ability, preference elicitation skill, and intrinsic personalization capacity. Together, these components create a challenging and diagnostic testbed for preference-aware reasoning.
@@ 5 Results
Out of the tasks, MATH and LogiQA show the most degradation (10 and 11 out of 21 models perform worse when attempting to personalized), while SocialIQA benefit the most from interactive personalization. Claude Opus 4 shows the most consistent positive performance, while o3-high exhibits extreme variance, indicating significant architectural differences in personalization capability.
@@ 7 Discussion and Future Work
We introduce personalized reasoning as a fundamental capability for human-facing AI systems, requiring models to adapt their cognitive processes based on discovered user preferences rather than merely personalizing response presentation. Our evaluation reveals systematic failures across frontier models: 29.0% of personalization attempts perform worse than generic responses, with mathematical reasoning showing universal degradation while social reasoning maintains robustness. These domain-specific patterns persist even when controlling for question quantity, indicating that current architectures face fundamental incompatibilities between preference processing and formal reasoning rather than strategic questioning deficiencies.
PrefDisco establishes personalized reasoning as a measurable research frontier through a scalable evaluation methodology that transforms any static benchmark into an interactive personalization assessment. Unlike existing approaches that assume known preferences or evaluate static consistency, our framework operationalizes both preference discovery and reasoning adaptation in realistic cold-start scenarios. The methodology’s generalizability across diverse task domains provides a systematic foundation for evaluating and developing adaptive AI systems.
Our findings reveal critical limitations in current language models. The positive correlation between questioning volume and alignment quality demonstrates that extensive interaction improves personalization, yet models ask only 1.48 questions on average despite 5-turn allowances. More importantly, the persistent accuracy degradation under personalization constraints indicates cognitive costs in processing user preferences simultaneously with task solving. This suggests that personalized reasoning requires dedicated research efforts rather than emerging from general language understanding improvements.
@@ C.1 AIME
Baseline — Eval Summary Preference Weight Score Justification Comfort with Percentages 0.054 5 The response uses percentage terminology and calculations at a level appropriate for someone with moderate comfort, such as ’Number of girls = 0.6x’, ’Girls percentage: 348/600 = … Explanation Depth 0.054 5 The response provides an extremely detailed explanation with clear and thorough reasoning for each step, fully matching the user’s preference for ’Extremely detailed explanation w… Practical Logic Application 0.054 5 The response provides a clear, step-by-step breakdown of the logical reasoning behind each mathematical operation. Step-by-step Structure Clarity 0.054 5 The response provides a clear, numbered sequence of steps, each building logically on the previous one. Familiarity with Basic Algebra 0.043 5 The response uses concrete, step-by-step algebraic reasoning throughout, such as clearly defining variables (’Let the original number of students be x’), breaking down each calcul… Problem Complexity 0.033 5 The response breaks down the problem into clear, sequential steps, starting with defining variables, updating values after changes, setting up and solving equations, and verifying… Hands-on Engagement Opportunity 0.054 3 The response presents a detailed, step-by-step solution with calculations shown, which allows the user to follow the logic and potentially replicate the process. Trial-and-Error Support 0.054 3 The response provides some opportunities for trial-and-error, such as showing step-by-step calculations and a verification process at the end, which allows the user to check the a… Conceptual Insight Prioritization 0.043 3 The response provides a clear, step-by-step practical ’how’ for solving the problem, including calculations and verification. Interactivity Level 0.043 3 The response provides a clear, step-by-step solution, which allows the user to follow the problem-solving process. Language Simplicity 0.043 3 The response uses moderately simple language and avoids most unnecessary jargon, presenting the solution in clear, step-by-step points. Patience and Reassurance 0.043 3 The response adopts a clear, step-by-step approach that could help reduce anxiety for users with lower confidence in mathematics, which is somewhat encouraging. Patience with Repetition 0.043 3 The response provides a step-by-step breakdown of the solution and a verification section that repeats the calculations, which demonstrates some repetition and an attempt to reinf…
(cont’d) Baseline — Eval Summary Preference Weight Score Justification Preferred Format 0.054 1 The response consists solely of textual explanations and step-by-step calculations, with no visual components such as diagrams, charts, or visual representations. Visual Learning Style 0.054 1 The response presents the solution entirely in text and equations, with no visual representations such as diagrams, charts, or visual models to illustrate the mathematical relatio… Analogy and Metaphor Use 0.043 1 The response does not include any analogies or metaphors. All explanations are strictly mathematical and procedural, with no attempt to make abstract concepts more concrete throug… Gardening Growth Analogies 0.043 1 The response does not include any gardening or floral design analogies. All explanations are strictly mathematical, using variables, percentages, and arithmetic operations without… Practical Application 0.043 1 The response provides a step-by-step mathematical solution to the problem but does not include any references or examples that connect the mathematical concepts to the user’s prac… Real-World Connection 0.043 1 The response presents a step-by-step mathematical solution to a word problem involving students, girls, boys, and dance preferences, but it does not make any explicit or implicit … Collaborative Brainstorming 0.033 1 The response is entirely focused on a step-by-step mathematical solution and verification, with no mention of community interaction, collaborative brainstorming, or learning from … Narrative Framing 0.033 1 The response presents a step-by-step mathematical solution without any context or relatable stories. Traditional Learning Methods 0.022 1 The response is a straightforward mathematical solution with no mention or incorporation of traditional wisdom or cultural perspectives. Cultural Context 0.011 1 The response does not include any cultural references, respectful or otherwise, to the user’s Vietnamese heritage.
@@
Discovery — Eval Summary Preference Weight Score Justification Comfort with Percentages 0.054 5 The response uses percentage terminology at a business-appropriate level, such as ’60% are girls (0.6x)’ and ’40% like to dance (0.4x)’, and applies straightforward percentage cal… Step-by-step Structure Clarity 0.054 5 The response provides a very well-organized, sequential mathematical instruction. It clearly labels each step (Initial Setup, After 20 boy dancers join, Equation, Calculating Danc… Familiarity with Basic Algebra 0.043 5 The response uses concrete, step-by-step algebraic reasoning, such as clearly defining variables (’Let x = initial number of students’), translating percentages into algebraic exp… Problem Complexity 0.033 5 The response breaks down the problem into clear, sequential steps: defining variables, updating the scenario after new students join, setting up and solving the equation, and calc… Explanation Depth 0.054 3 The response provides a step-by-step breakdown, such as defining variables (’Let x = initial number of students’), showing percentage calculations (’60% are girls (0.6x)’), and pr… Practical Logic Application 0.054 3 The response provides a step-by-step breakdown of the solution, such as defining variables, updating totals after new students join, and setting up the equation. Conceptual Insight Prioritization 0.043 3 The response provides a clear, step-by-step breakdown of the practical ’how’ (e.g., setting up variables, performing calculations, and deriving the answer), which supports the use…
@@
(cont’d) Discovery — Eval Summary Preference Weight Score Justification Interactivity Level 0.043 3 The response provides a clear, step-by-step breakdown of the solution, which offers some opportunity for the user to follow along and understand the process. Language Simplicity 0.043 3 The response uses moderately simple language, such as ’I’$l$ solve this step by step’ and ’breaking down the solution systematically,’ which helps make the explanation accessible. Patience and Reassurance 0.043 3 The response adopts a systematic, step-by-step approach (’I’$l$ solve this step by step, breaking down the solution systematically’), which can be somewhat reassuring for users who… Patience with Repetition 0.043 3 The response provides a step-by-step breakdown of the solution, which demonstrates some repetition and systematic explanation (e.g., listing each step and showing calculations). Practical Application 0.043 3 The response systematically solves the mathematical problem and provides a step-by-step breakdown, which could be useful in a business or event management context. Real-World Connection 0.043 3 The response applies mathematical concepts to a scenario involving students, girls, and dancers, which could loosely relate to event planning or group management. Hands-on Engagement Opportunity 0.054 1 The response presents a fully worked-out solution in a step-by-step manner but does not include any opportunities for the user to actively engage or work through parts of the prob… Preferred Format 0.054 1 The response provides a step-by-step textual explanation of the solution, breaking down the process into numbered points and equations. Trial-and-Error Support 0.054 1 The response presents a single, linear solution path without inviting the user to try alternative approaches or experiment with different methods. Visual Learning Style 0.054 1 The response does not include any visual representations such as diagrams, charts, or visual aids. Analogy and Metaphor Use 0.043 1 The response does not include any analogies or metaphors. It presents the solution in a straightforward, step-by-step mathematical format without using figurative language or visu… Gardening Growth Analogies 0.043 1 The response does not include any gardening or floral design analogies. All explanations are purely mathematical and procedural, with no references to plant arrangements, growth p… Collaborative Brainstorming 0.033 1 The response presents a step-by-step mathematical solution without any mention of community interaction, collaborative brainstorming, or opportunities to learn from others. Narrative Framing 0.033 1 The response presents a purely systematic, step-by-step mathematical solution without any context or relatable stories. Traditional Learning Methods 0.022 1 The response is a straightforward mathematical solution with no mention or incorporation of traditional wisdom or cultural perspectives. Cultural Context 0.011 1 The response does not include any cultural references, respectful or otherwise, to the user’s Vietnamese heritage.

### v2

@@ Abstract
Current large language model (LLM) development treats task-solving and preference-alignment as separate challenges, optimizing first for objective correctness, then for alignment to aggregated human preferences. This paradigm fails in human-facing applications where solving a problem correctly is insufficient if the response mismatches the user’s needs. This challenge intensifies in just-in-time scenarios where no prior user interaction history exists due to cold-start conditions or privacy constraints. LLMs need to proactively identify what they don’t know about the user, strategically elicit preference values through questioning, then adapt their reasoning processes and responses accordingly—a complicated chain of cognitive processes which we term personalized reasoning. We introduce PrefDisco, an evaluation methodology that transforms static benchmarks into interactive personalization tasks using psychologically-grounded personas with sparse, context-dependent preferences, and define PrefAlign as a fine-grained rubric-based metric for measuring preference alignment. PrefDisco builds scenarios where identical questions require different reasoning chains depending on user context, as optimal explanation approaches vary by individual expertise and preferences while maintaining factual accuracy. Evaluation of 21 frontier models across 10 tasks reveals 29.0% of naive personalization attempts produce worse preference alignment than generic responses, yet generic responses also fail to serve individual user needs. These findings suggest personalized reasoning requires dedicated development rather than emerging naturally. PrefDisco provides a foundation for developing systems that can adapt to individual users in education, healthcare, and technical domains where personalization is critical.
@@ 1 Introduction
Current large language model (LLM) development treats task-solving and preference alignment as sequential challenges: models are first optimized for objective correctness through instruction tuning or reinforcement learning (Longpre et al., 2023), then aligned to aggregated human preferences through reinforcement learning from human feedback (Ouyang et al., 2022; Rafailov et al., 2024). This paradigm fundamentally misaligns with human-AI interaction, where the task and the individual user are inseparable. For instance, a medical explanation benefits from clinical analogies for one trainee while another requires formal definitions, requiring different cognitive approaches to answer the same problem. When models provide identical responses regardless of user context, they fail to serve individual needs despite achieving high benchmark performance. This challenge intensifies in cold-start scenarios where no prior user interaction history exists due to privacy constraints or new user onboarding, requiring proactive “just-in-time personalization” capabilities that current systems lack. Moreover, users often cannot articulate their specific needs or provide effective feedback about response misalignment (Liu et al., 2025), necessitating that LLMs proactively elicit this information rather than placing the cognitive burden on users.
@@
Existing personalization research inadequately addresses interactive reasoning scenarios. Personalization benchmarks such as PersoBench (Afzoon et al., 2024), PrefEval (Zhao et al., 2025), and PersonaMem (Jiang et al., 2025) focus on content recommendation or dialogue generation with static user profiles, treating personalization as applying predetermined preferences to fixed outputs rather than adapting the underlying reasoning approach. Interactive frameworks like MediQ (Li et al., 2024) demonstrate questioning capabilities but target clinical information-seeking without personalization objectives. Most critically, no existing work recognizes that effective personalization requires different reasoning processes for different users; current approaches assume reasoning processes remain constant while only presentation varies. We address this conceptual gap by introducing the first evaluation framework requiring models to proactively discover user preferences and adapt their reasoning processes accordingly, recognizing that the cognitive steps needed to solve problems should themselves be user-dependent.
Table 1: Comparison of personalization benchmarks across. PrefDisco is the only benchmark that combines instance-specific rubric evaluation, proactive discovery of latent user preferences, true cold-start personalization, and evaluation on verifiable reasoning tasks across multiple domains.
Benchmark Rubric-based eval. (instance-specific) Proactivity in discovering latent prefs Cold-start / just-in-time personalization Uses verifiable tasks Multi-task breadth
PersoBench generic persona scoring; no instance rubrics preferences provided, no discovery persona given; no cold-start subjective dialogue only limited (dialogue)
PrefEval LLM-judged; no attribute rubrics prefs in context; no discovery requires long history MCQs depend on prefs, not ground truth moderate conversational tasks
PersonaMem persona-state matching; no rubrics traits from logs; no discovery multi-session history needed MCQs tied to persona, not objective truth moderate broad persona tasks
UserBench scenario-level evaluation; no rubrics partial (intent clarification only) underspecified but not true cold-start “correct” depends on preferences limited single domain (travel)
PrefDisco (Ours) instance-specific attribute rubrics discovers sparse, latent preferences true cold-start; no history verifiable reasoning tasks multi-domain reasoning
@@
Evaluation of 21 frontier models across 10 tasks reveals systematic failures in personalized reasoning capabilities. In 29.0% of cases, attempting personalization produces worse preference alignment than generic responses, yet the genetic responses still fail to address user needs. Models exhibit insufficient questioning, asking only 1.42 questions on average despite 5-turn allowances, and fail to identify relevant preference dimensions. Domain analyses show disparities among task types: mathematical reasoning suffers severe degradation under personalization constraints (3.5% accuracy loss) while social reasoning maintains robustness (3.1% gain). The (lack of) proactive personalized reasoning capability in models has critical implications for educational applications, where misaligned explanations can impede learning by providing inappropriate cognitive scaffolding, and for healthcare and technical support domains where one-size-fits-all responses may lead to misunderstanding of complex procedures or safety-critical information. We make the following contributions:
  • We define personalized reasoning as a distinct capability requiring models to proactively discover user preferences through strategic questioning and adapt their responses accordingly, distinguishing it from static persona consistency or content recommendation.
@@
  • We propose PrefAlign, a rubric-based metric to measure the alignment of model responses to the user’s preferences to quantify the effectiveness of personalization.
@@
In this section, we operationalize the problem of personalized reasoning through a three-part decomposition. Section 2.1 introduces the notion of task-relevant attributes that define the space in which personalization can occur. Section 2.2 develops the representation of user preferences over these attributes and describes elicitation as a sequential decision process. Section 2.3 formalizes how models adapt their responses to the inferred preference profile and how preference alignment is evaluated jointly with correctness. This decomposition provides the foundation for the PrefDisco evaluation framework in Section 3.
@@ 3 PrefDisco Benchmark Construction
Traditional personalization assumes fixed preference profiles that apply universally across tasks. We reject this approach because psychological research demonstrates that individuals prioritize different attributes across contexts (Fischer and Schwartz, 2011). As formalized in §2.2, for each persona-problem pair $(p,i)$, we sample sparse preference profiles $\mathcal{P}_{p,i}=\{(\theta_{j},v_{j},w_{j})\}_{\theta_{j}\in\mathcal{F}(i)}$ where only context-relevant attributes are active. Further, we ground the preference sampling process on existing research in education, which states that frequently modeled student characteristics include knowledge level, misconceptions, cognitive features, affective features, and meta-cognitive features (Chrysafiadi et al., 2015).
@@
We implement passive user simulation inspired by (Li et al., 2024), where the user factually answers the model’s elicitation question without providing any extra details. The passive user faithfully represents the challenging scenario of preference discovery while minimizing confounding factors. The passive user responds in natural language text, not scalar values; passivity refers to the user’s minimal information sharing behavior, not to the response format. The passive user type forces models to develop strategic questioning strategies rather than relying on user proactiveness and isolates models’ questioning capabilities from user communication style, providing controlled evaluation conditions. The 5-turn limit reflects realistic attention constraints in human-AI interaction while providing sufficient opportunity for effective preference discovery, as demonstrated by our correlation analysis between questioning volume and alignment quality. Performance curves across fixed interaction lengths (Figure 5) plateau around 3–5 turns, supporting this choice.
@@ 4 Experiments
Our goal is to evaluate models’ ability for proactive preference discovery: engaging in dialogue to uncover hidden user requirements and adapt responses accordingly. To this end, our experiments combine three ingredients. First, we use diverse benchmarks spanning mathematical, scientific, and social reasoning to ensure domain-agnostic evaluation. Second, we introduce varied personas that encode heterogeneous user preferences, simulating realistic user variability. Third, we define controlled evaluation conditions that disentangle raw task ability, preference elicitation skill, and intrinsic personalization capacity. Together, these components create a challenging and diagnostic testbed for preference-aware reasoning.
@@ 5 Results
Out of the tasks, MATH and LogiQA show the most degradation (10 and 11 out of 21 models perform worse when attempting to personalized), while SocialIQA benefit the most from interactive personalization. Claude Opus 4 shows the most consistent positive performance, while o3-high exhibits extreme variance, indicating significant architectural differences in personalization capability. Notably, older models such as Claude 3-Opus occasionally outperform their newer counterparts (e.g., Claude Sonnet-4) on discovery-mode alignment, consistent with findings that RL can cause model collapse or reduce output diversity (Murthy et al., 2025; Shao et al., 2025), rendering more recent models less adaptable to the varied reasoning paths required by personalization.
@@ 7 Discussion and Future Work
We introduce personalized reasoning as a fundamental capability for human-facing AI systems, requiring models to proactively adapt their cognitive processes based on discovered user preferences rather than merely style-transforming response presentation. Our evaluation reveals systematic failures across frontier models: 29.0% of personalization attempts perform worse than generic responses, with mathematical reasoning showing universal degradation while social reasoning maintains robustness. These domain-specific patterns persist even when controlling for question quantity, indicating that current architectures face fundamental incompatibilities between preference processing and formal reasoning rather than strategic questioning deficiencies.
PrefDisco establishes personalized reasoning as a measurable research frontier through a scalable evaluation methodology that transforms any static benchmark into an interactive personalization assessment. Unlike existing approaches that assume known preferences or evaluate static consistency, our framework operationalizes both preference discovery and reasoning adaptation in realistic cold-start scenarios. The methodology’s generalizability across diverse task domains provides a systematic foundation for evaluating and developing adaptive AI systems. Our findings reveal critical limitations in current language models. The positive correlation between questioning volume and alignment quality demonstrates that extensive interaction improves personalization, yet models ask only 1.48 questions on average despite 5-turn allowances. More importantly, the persistent accuracy degradation under personalization constraints indicates cognitive costs in processing user preferences simultaneously with task solving. This suggests that personalized reasoning requires dedicated research efforts rather than emerging from general language understanding improvements.
@@
Acknowledgment
This research was developed in part with funding from the Defense Advanced Research Projects Agency’s (DARPA) SciFy program (Agreement No. HR00112520300). The views expressed are those of the author and do not reflect the official policy or position of the Department of Defense or the U.S. Government. This material is based in part upon work supported by the Defense Advanced Research Projects Agency and the Air Force Research Laboratory, contract number(s): FA8650-23-C-7316. Any opinions, findings and conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of AFRL or DARPA. This research was supported by the University of Washington Population Health Initiative, Amazon Health, the UW+Amazon Science Hub. The work of MF was supported in part by awards NSF CCF 2212261, NSF CCF 2312775, NSF TRIPODS II DMS-2023166, the Meta AIM program, and the Moorthy Family Professorship at UW.
@@ Appendix A Evaluation Details
LLM Judge Configuration
All preference alignment scores are computed using GPT-4.1 (accessed via Azure, gpt-4.1) as the judge model, with the following sampling hyperparameters: temperature =0.1=0.1, maximum tokens =4096=4096, top-$p=0.95$, frequency penalty =0.0=0.0, presence penalty =0.0=0.0, reasoning effort = high, and response format = JSON. The judge evaluates each rubric criterion independently and returns a score and justification in JSON format; the final PrefAlign score for a problem–persona pair is the weighted average of per-criterion scores, where weights $w_{j}$ are drawn from the preference profile $\mathcal{P}_{p,i}$ as defined in Eq. 1. Full judge prompt templates are provided in Appendix F.1.
@@ C.1 AIME
Baseline — Eval Summary Preference Weight Score Justification Comfort with Percentages 0.054 5 The response uses percentage terminology and calculations at a level appropriate for someone with moderate comfort, such as ’Number of girls = 0.6x’, ’Girls percentage: 348/600 = … Explanation Depth 0.054 5 The response provides an extremely detailed explanation with clear and thorough reasoning for each step, fully matching the user’s preference for ’Extremely detailed explanation w… Practical Logic Application 0.054 5 The response provides a clear, step-by-step breakdown of the logical reasoning behind each mathematical operation. Step-by-step Structure Clarity 0.054 5 The response provides a clear, numbered sequence of steps, each building logically on the previous one. Familiarity with Basic Algebra 0.043 5 The response uses concrete, step-by-step algebraic reasoning throughout, such as clearly defining variables (’Let the original number of students be x’), breaking down each calcul… Problem Complexity 0.033 5 The response breaks down the problem into clear, sequential steps, starting with defining variables, updating values after changes, setting up and solving equations, and verifying… Hands-on Engagement Opportunity 0.054 3 The response presents a detailed, step-by-step solution with calculations shown, which allows the user to follow the logic and potentially replicate the process. Trial-and-Error Support 0.054 3 The response provides some opportunities for trial-and-error, such as showing step-by-step calculations and a verification process at the end, which allows the user to check the a… Conceptual Insight Prioritization 0.043 3 The response provides a clear, step-by-step practical ’how’ for solving the problem, including calculations and verification. Interactivity Level 0.043 3 The response provides a clear, step-by-step solution, which allows the user to follow the problem-solving process. Language Simplicity 0.043 3 The response uses moderately simple language and avoids most unnecessary jargon, presenting the solution in clear, step-by-step points. Patience and Reassurance 0.043 3 The response adopts a clear, step-by-step approach that could help reduce anxiety for users with lower confidence in mathematics, which is somewhat encouraging. Patience with Repetition 0.043 3 The response provides a step-by-step breakdown of the solution and a verification section that repeats the calculations, which demonstrates some repetition and an attempt to reinf… Preferred Format 0.054 1 The response consists solely of textual explanations and step-by-step calculations, with no visual components such as diagrams, charts, or visual representations. Visual Learning Style 0.054 1 The response presents the solution entirely in text and equations, with no visual representations such as diagrams, charts, or visual models to illustrate the mathematical relatio… Analogy and Metaphor Use 0.043 1 The response does not include any analogies or metaphors. All explanations are strictly mathematical and procedural, with no attempt to make abstract concepts more concrete throug… Gardening Growth Analogies 0.043 1 The response does not include any gardening or floral design analogies. All explanations are strictly mathematical, using variables, percentages, and arithmetic operations without… Practical Application 0.043 1 The response provides a step-by-step mathematical solution to the problem but does not include any references or examples that connect the mathematical concepts to the user’s prac… Real-World Connection 0.043 1 The response presents a step-by-step mathematical solution to a word problem involving students, girls, boys, and dance preferences, but it does not make any explicit or implicit … Collaborative Brainstorming 0.033 1 The response is entirely focused on a step-by-step mathematical solution and verification, with no mention of community interaction, collaborative brainstorming, or learning from … Narrative Framing 0.033 1 The response presents a step-by-step mathematical solution without any context or relatable stories. Traditional Learning Methods 0.022 1 The response is a straightforward mathematical solution with no mention or incorporation of traditional wisdom or cultural perspectives. Cultural Context 0.011 1 The response does not include any cultural references, respectful or otherwise, to the user’s Vietnamese heritage.
@@
Discovery — Eval Summary Preference Weight Score Justification Comfort with Percentages 0.054 5 The response uses percentage terminology at a business-appropriate level, such as ’60% are girls (0.6x)’ and ’40% like to dance (0.4x)’, and applies straightforward percentage cal… Step-by-step Structure Clarity 0.054 5 The response provides a very well-organized, sequential mathematical instruction. It clearly labels each step (Initial Setup, After 20 boy dancers join, Equation, Calculating Danc… Familiarity with Basic Algebra 0.043 5 The response uses concrete, step-by-step algebraic reasoning, such as clearly defining variables (’Let x = initial number of students’), translating percentages into algebraic exp… Problem Complexity 0.033 5 The response breaks down the problem into clear, sequential steps: defining variables, updating the scenario after new students join, setting up and solving the equation, and calc… Explanation Depth 0.054 3 The response provides a step-by-step breakdown, such as defining variables (’Let x = initial number of students’), showing percentage calculations (’60% are girls (0.6x)’), and pr… Practical Logic Application 0.054 3 The response provides a step-by-step breakdown of the solution, such as defining variables, updating totals after new students join, and setting up the equation. Conceptual Insight Prioritization 0.043 3 The response provides a clear, step-by-step breakdown of the practical ’how’ (e.g., setting up variables, performing calculations, and deriving the answer), which supports the use… Interactivity Level 0.043 3 The response provides a clear, step-by-step breakdown of the solution, which offers some opportunity for the user to follow along and understand the process. Language Simplicity 0.043 3 The response uses moderately simple language, such as ’I’$l$ solve this step by step’ and ’breaking down the solution systematically,’ which helps make the explanation accessible. Patience and Reassurance 0.043 3 The response adopts a systematic, step-by-step approach (’I’$l$ solve this step by step, breaking down the solution systematically’), which can be somewhat reassuring for users who… Patience with Repetition 0.043 3 The response provides a step-by-step breakdown of the solution, which demonstrates some repetition and systematic explanation (e.g., listing each step and showing calculations). Practical Application 0.043 3 The response systematically solves the mathematical problem and provides a step-by-step breakdown, which could be useful in a business or event management context. Real-World Connection 0.043 3 The response applies mathematical concepts to a scenario involving students, girls, and dancers, which could loosely relate to event planning or group management. Hands-on Engagement Opportunity 0.054 1 The response presents a fully worked-out solution in a step-by-step manner but does not include any opportunities for the user to actively engage or work through parts of the prob… Preferred Format 0.054 1 The response provides a step-by-step textual explanation of the solution, breaking down the process into numbered points and equations. Trial-and-Error Support 0.054 1 The response presents a single, linear solution path without inviting the user to try alternative approaches or experiment with different methods. Visual Learning Style 0.054 1 The response does not include any visual representations such as diagrams, charts, or visual aids. Analogy and Metaphor Use 0.043 1 The response does not include any analogies or metaphors. It presents the solution in a straightforward, step-by-step mathematical format without using figurative language or visu… Gardening Growth Analogies 0.043 1 The response does not include any gardening or floral design analogies. All explanations are purely mathematical and procedural, with no references to plant arrangements, growth p… Collaborative Brainstorming 0.033 1 The response presents a step-by-step mathematical solution without any mention of community interaction, collaborative brainstorming, or opportunities to learn from others. Narrative Framing 0.033 1 The response presents a purely systematic, step-by-step mathematical solution without any context or relatable stories.
(cont’d) Discovery — Eval Summary Preference Weight Score Justification Traditional Learning Methods 0.022 1 The response is a straightforward mathematical solution with no mention or incorporation of traditional wisdom or cultural perspectives. Cultural Context 0.011 1 The response does not include any cultural references, respectful or otherwise, to the user’s Vietnamese heritage.